OpenStudy (anonymous):

The container that holds the water for the football team is 1/2 full. After pouring out 2 gallons of water, it is 1/3 full. How many gallons can the container hold?

jimthompson5910 (jim_thompson5910):

x = total number of gallons the container holds (when it is filled up to the top) currently it has (1/2)*x = x/2 gallons of water pour our 2 gallons and we end up with (x/2) - 2 since it's now 1/3 full, it has x/3 gallons of water left so this is the equation (x/2) - 2 = x/3

jimthompson5910 (jim_thompson5910):

to clear out the fractions, multiply both sides by 6 (x/2) - 2 = x/3 6*[(x/2) - 2] = 6*(x/3) 6*(x/2) + 6*(-2) = 6*(x/3) 6x/2 - 12 = 6x/3 3x - 12 = 2x I'll let you finish up

OpenStudy (anonymous):

12!

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Can you help me with more? I'd be so appreciative and solve all of it